In a random sample of size 49, 29 people say they generally get less than 8 hours of sleep a night. Calculate the point estimates p^ and q^
Answer #1

Here, we have given that,

n = total number of people = 49

x= number of people who say generally they get less than 8 hours of sleep out of total 49 people=29

we know that the \hat p and \hat q best point estimate of the population.

Now, we want to find \hat p and \hat q

\hat p=sample proportion = probability of success

=\frac{ x} {n}

=\frac{29}{49}

= 0.59

Now,

\hat q = probability of failure

= 1- \hat p

=1 - 0.59

= 0.41

Here, we get the

\hat p= 0.59 and \hat q = 0.41